OpenStudy (anonymous):

HELP PLEASE!!! I need two numbers that multiply to give you -20 and add up to give you -5. factoring trinomials

OpenStudy (anonymous):

can you write your original equation something doesn't fit

OpenStudy (campbell_st):

whats the coefficient of x^2

OpenStudy (campbell_st):

because I think you need to solve using general quadratic formula...or completing the square

OpenStudy (anonymous):

a^2-5a-20

OpenStudy (anonymous):

and its factoring trinomials

OpenStudy (anonymous):

\[a^2-5a-20\]

OpenStudy (anonymous):

yup

OpenStudy (campbell_st):

the factors of -20 are 1 x 20, 2 x 10 or 4 x 5 you can't get five from the factors with the the larger factor being negative...

OpenStudy (anonymous):

pretty sure @campbell_st is correct using the general quadratic formula otherwise like he just said its "not possible".

OpenStudy (anonymous):

then the problem is prime

OpenStudy (campbell_st):

using the general quadratic formula the will give you factors of\[x = \frac{5 + \sqrt{105}}{2} ======== and===== x = \frac{5 - \sqrt{105}}{2}\]

OpenStudy (anonymous):

im in 9th grade this is algebra 1

OpenStudy (anonymous):

lol oh

OpenStudy (anonymous):

yah so i have no idea wat tht is

OpenStudy (campbell_st):

well just tell you teacher there is a typing error in the question...

OpenStudy (anonymous):

:)^^^^

OpenStudy (anonymous):

or say you did it using the quadratic formula ;)

OpenStudy (anonymous):

or the problem is prime

OpenStudy (anonymous):

oh well i think its prime

OpenStudy (campbell_st):

it means it can't be factored... so yes its prime

OpenStudy (anonymous):

yes if thats the case then its prime

OpenStudy (anonymous):

ok then i was right thanks guys for all ur help! :D